An average computer mouse inspector can inspect 45 mice per hour with a population standard deviation of 14 mice per hour. The 44 computer mice inspectors at a particular factory can only inspect 38 mice per hour. Does the company have reason to believe that these inspectors are slower than average at α = 0.10?
Show work, thanks

Answer by stanbon(75887)

An average computer mouse inspector can inspect 45 mice per hour with a population standard deviation of 14 mice per hour. The 44 computer mice inspectors at a particular factory can only inspect 38 mice per hour. Does the company have reason to believe that these inspectors are slower than average at α = 0.10?
----------
Ho: u = 45
Ha: u < 45
-------
z(38) = (38-45)/[14/sqrt(44)] = -3.3166
-----
p-value = P(z < -3.3166) = 0.00046
-----
Conclusion: Since the p-value is less than 10%,reject Ho at the 10% level
of significance.
----
Cheers,
Stan H.
===========